Find all the real number roots of the equation.
log2(x+7)-log2(x-3)=3

Answer by lwsshak3(11628)

Find all the real number roots of the equation.
log2(x+7)-log2(x-3)=3
log2[(x+7)/(x-3)]=3
convert to exponential form
(x+7)/(x-3)=2^3=8
x+7=8x-24
7x=31
x=31/7=4.4286..